OpenStudy (anonymous):

A certain forest covers an area of 3200km^2. Suppose that each year this area decreases by 3.25%. What will the area be after 7 years?

OpenStudy (joyeezy):

Area after 12 years = 4300(1 - 0.07)^12 = 4300(0.93)^12 = 1799.96 km^2